# The decomposition of primes in nonabelian extensions of Heisenberg type and an analogue of Euler’s criterion

### v2

@@ 1. Introduction
Since the class field theory was founded in the early 20th century [CO], whose scope is limited to abelian extensions, it has been a fascinating problem to find laws which govern the prime decomposition in nonabelian extensions of global fields. For quadratic extensions, the decomposition pattern is encoded in the Legendre symbol and the desired law corresponds to that of quadratic reciprocity. Combined with Euler’s criterion, the law says
@@
for distinct odd primes $p$ and $q$ such that $4$ divides p1p-1. In the language of class field theory, it relates the Artin symbols at $q$ in two abelian extensions $\mathbb{Q}(\sqrt{p})$ and $\mathbb{Q}(\zeta_{p})$ of $\mathbb{Q}$, where $\zeta_{m}$ denotes a primitive $m$-th root of unity for $m\geq 1$.
In this paper, we establish an analogue of (1) for certain nonabelian extensions of function fields. The particular extensions we consider are those which arise naturally from attempts to generalize Legendre symbols using Massey products. We mention that Massey products in Galois cohomology have received much attention in their own right, including some recent works [Harpaz-Wittenberg, Maire-et-al]. For a specific context regarding the multilinear extension of the Legendre symbol, we refer interested readers to [MO] for a survey and to closely related works [AF, KU, AM3, AMN]. Our field extensions first appeared in [HM], which is recalled below.
@@
Temporarily, denote by $k$ an arbitrary field containing a primitive $\ell$-th root of unity $\zeta_{\ell}$. In [HM], the authors introduced a Galois extension $\mathcal{R}^{(\ell)}$ of the function field $k(t)$ with Galois group $\mathcal{H}(\mathbb{F}_{\ell})$, which we call a mod $\ell$ Heisenberg extension. It is defined as $\mathcal{R}^{(\ell)}\coloneq k(t)(t^{1/\ell},(1-t)^{1/\ell},\epsilon_{\ell}(t)^{1/\ell})$ where $\epsilon_{\ell}(t)\coloneq\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i}t^{1/\ell})^{i}$. They utilized $\mathcal{R}^{(\ell)}$ to establish a relationship between two arithmetic invariants: the mod $\ell$ Milnor invariants $\mu_{\ell}(g;I)$ of Galois elements defined in [KMT] and the mod $\ell$ Milnor invariants $\mu_{\ell}(J)$ of primes introduced in [MO]. Also, they discovered analogies between $\mathcal{R}^{(\ell)}$ and dilogarithmic functions. From this viewpoint, the authors of [NA, SH] have explored related structures involving $\ell$-adic Galois polylogarithms and triple $\ell$-th power symbols. We remark that these triple power symbols are regarded as multilinear generalizations of Legendre symbols.
For our purpose, we will take $k=\mathbb{F}_{p}$ for a prime $p$. Since we need a primitive $\ell$-th root of unity, we always assume $\ell\mid p-1$. Our main result concerns the splitting of the prime ideal (ta)(t-a) for $a\in\mathbb{F}_{p}-\{0,1\}$ in the mod $\ell$ Heisenberg extension $\mathcal{R}^{(\ell)}/\mathbb{F}_{p}(t)$. To state it, we introduce certain polynomials that will play the role of $x^{(p-1)/2}$ for $\mathbb{Q}(\sqrt{p})$. Put $\epsilon_{\ell,n}(t)=\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i+n}t^{1/\ell})^{i}$. Note that $\epsilon_{\ell,0}(t)$ is literally $\epsilon_{\ell}(t)$ which is a generator of $\mathcal{R}^{(\ell)}$. Define the polynomial
(2) $A_{\ell}(x):=\begin{cases}\frac{1}{2}\left((1-\sqrt{x})^{\frac{p-1}{2}}+(1+\sqrt{x})^{\frac{p-1}{2}}\right)&\text{ if $\ell=2$,}\\ \frac{1}{\ell}\left(\sum_{j=0}^{\ell-1}\epsilon_{\ell,j}(x)^{\frac{p-1}{l}}\right)&\text{ if $\ell>2$.}\end{cases}$
Indeed, $A_{2}(x)$ is a polynomial in $x$ since it is invariant under the substitution $\sqrt{x}\mapsto-\sqrt{x}$. Similarly, one can show that $A_{\ell}(x)$ is also a polynomial in $x$ using the substitution $t^{1/\ell}\mapsto\zeta_{\ell}t^{1/\ell}$.
@@
In the case $\ell=2$, we have: {restatable}theoremspliteven Let $p$ be an odd prime and $a\in\mathbb{F}_{p}$ which $a\neq 0,1,\frac{1}{2}$. Let $n_{a}$ be the number of primes above (ta)(t-a) in $\mathcal{R}^{(2)}/\mathbb{F}_{p}(t)$. Then, $n_{a}$ is determined by the value of $A_{2}(a)$;
@@
The reason we separate the case $\ell=2$ is reflected in the group structure of $\mathcal{H}(\mathbb{F}_{\ell})$ which has an element of order $\ell^{2}$ precisely when $\ell=2$. As an another feature of the case $\ell=2$, we identify the morphism between algebraic curves underlying $\mathcal{R}^{(\ell)}/k(t)$ which guide the necessary computation. For a general $\ell$, we do not have such geometric interpretation and rely on algebraic manipulations. Despite of the difference, a key ingredient common to all cases is an explicit integral basis of $\mathcal{R}^{(\ell)}$. The construction of such a basis seems new for $\ell\geq 3$ and it is the most technical part of the paper.
@@ 2.1. Frobenius conjugacy classes
    Let $\mathfrak{q}$ be a prime ideal of $B$ lying over $\mathfrak{p}$, and $f$ be the order of $\operatorname{Frob}_{\mathfrak{p},\mathfrak{q}}\in\operatorname{Gal}(L/K)$. Then, the number of prime ideals of $\mathcal{O}_{L}$ lying above $\mathfrak{p}$ equals to $\frac{1}{f}[L:K]$.
@@ 2.3. Mod $\ell$ Heisenberg Extension
Using the above notation, $\mathcal{R}^{(\ell)}/k(t)$ is a Galois extension such that $\operatorname{Gal}(\mathcal{R}^{(\ell)}/k(t))$ is isomorphic to $\mathcal{H}(\mathbb{F}_{\ell})$. Moreover, it is unramified outside $\{\infty,0,1\}$.
@@
induces the isomorphism $\operatorname{Gal}(\mathcal{R}^{(\ell)}/k(t))\cong H(\mathbb{F}_{\ell}).$
@@ 3. $\ell=2$
Since the Frobenius conjugacy class that corresponds to (ta)(t-a) has order of at most $4$, the number of prime ideals above (ta)(t-a) is at least $2$ by Proposition 1. ∎
@@
Here, recall that $\mathcal{C}^{(2)}$ is the Fermat plane curve $X^{2}+Y^{2}=Z^{2}$ in $\mathbb{P}^{2}$ corresponding to $\mathcal{K}^{(2)}$. Since $\genfrac{(}{)}{}{}{a}{p}=\genfrac{(}{)}{}{}{1-a}{p}=1$, the splitting of (ta)(t-a) in $\mathcal{K}^{(2)}$ can be represented by four closed points $(\pm x:\pm y:1)$ where $x,y\in\mathbb{F}_{p}$ that satisfies $x^{2}=a,y^{2}=1-a$.
@@
$\genfrac{(}{)}{}{}{1-x}{p}=\frac{1}{2}\left(\genfrac{(}{)}{}{}{1-x}{p}+\genfrac{(}{)}{}{}{1+x}{p}\right)=\frac{1}{2}\left((1-x)^{\frac{p-1}{2}}+(1+x)^{\frac{1-p}{2}}\right).$
Recall from (2) that $A_{2}(a)=\frac{1}{2}\left((1-\sqrt{a})^{\frac{p-1}{2}}+(1+\sqrt{a})^{\frac{p-1}{2}}\right)$. So, if $\genfrac{(}{)}{}{}{1-x}{p}=1$, the value of the above polynomial is $1$, and if $\genfrac{(}{)}{}{}{1-x}{p}=-1$, the value of the above polynomial is 1-1. Finally, since $x^{2}=a$ in $\mathbb{F}_{p}$ and $\frac{1}{2}\left((1-x)^{\frac{p-1}{2}}+(1+x)^{\frac{1-p}{2}}\right)$ only has even degree terms, we obtain our desired result. ∎
Now, we show $A_{2}(a)$ determines other quadratic residue symbols.
@@
Note that $A_{2}(a)=x_{\frac{p-1}{2}}$. Then, $x_{0}=1,x_{1}=1$ and $x_{n}$ satisfies the recursive relation $x_{n+1}=2x_{n}+(a-1)x_{n-1}$. Also, from the definition it is straightforward to see that $x_{2n}=2x_{n}^{2}-(1-a)^{n}$ holds. Now, for $a\in\mathbb{F}_{p}$,
$x_{p}=\frac{1}{2}((1-\sqrt{a})^{p}+(1+\sqrt{a})^{p}=\frac{1}{2}\sum_{i=0}^{p}(1+(-1)^{i})\binom{p}{i}(\sqrt{a})^{i}=1$
where the first equality is the definition, the second equality is the binomial expansion, and the third equality holds because all but the first term vanishes. In the last equality, we rely on $p$ being odd when $\ell=2$. A similar argument shows that $x_{p+1}=\frac{1}{2}((1-\sqrt{a})^{p+1}+(1+\sqrt{a})^{p+1})=1+a^{\frac{p+1}{2}}=1+a^{\frac{p-1}{2}}\cdot a$.
Now, consider the case $a^{\frac{p-1}{2}}=1$. In this case, $x_{p+1}=1+a$ and we get $x_{p-1}=1$ from the recursive formula for $x_{n}$. Then,
$x_{\frac{p-1}{2}}^{2}=\frac{1}{2}\left(1+(1-a)^{\frac{p-1}{2}}\right).$
@@
In the other case $a^{\frac{p-1}{2}}=-1$, we have $a_{p+1}=1-a$. Again from the recursive formula, we get $a_{p-1}=\frac{a+1}{1-a}$. Then,
$x_{\frac{p-1}{2}}^{2}=\frac{1}{2}\left(\frac{a+1}{1-a}+(1-a)^{\frac{p-1}{2}}\right).$
If $(1-a)^{\frac{p-1}{2}}=1$, $x_{\frac{p-1}{2}}^{2}=\frac{1}{1-a}$ which is never $0,1$. If $(1-a)^{\frac{p-1}{2}}=-1$, $x_{\frac{p-1}{2}}^{2}=\frac{a}{1-a}$ which is not $0,1,\frac{1}{1-a}$ except the case $a=\frac{1}{2}$. Combining these results, we proved the desired result. ∎
@@
From this calculation, we observe that for a representative $(a_{ij})\in\mathcal{H}(\mathbb{F}_{2})$ of a conjugacy class, if $a_{12}=0,a_{21}=1$ or $a_{12}=1,a_{21}=0$ holds, $(a_{ij})$ has order 2. Otherwise, if $a_{12}=a_{21}=1$, $(a_{ij})$ has order 4.
Recall that as an element of the Galois group $\operatorname{Gal}(\mathcal{R}^{(2)}/k(t))\cong\mathcal{H}(\mathbb{F}_{2})$, $a_{12}$ is determined from $\sqrt{t}\mapsto\pm\sqrt{t}$, $a_{21}$ from $\sqrt{1-t}\mapsto\pm\sqrt{1-t}$, $a_{13}$ from $\sqrt{1-\sqrt{t}}\mapsto\pm\sqrt{1-a_{12}\sqrt{t}}$. When we restrict $\operatorname{Frob}_{(t-a),\mathcal{R}^{(2)}}$ to $K_{1}=\mathbb{F}_{p}(t)(\sqrt{t})$, we get $\operatorname{Frob}_{(t-a),K_{1}}$. Since $K_{1}/\mathbb{F}_{p}(t)$ is abelian, its induced morphism to $\mathcal{O}_{K_{1}}/(t-a)\to\mathcal{O}_{K_{1}}/(t-a)$ is given by $x\mapsto x^{p}$. Hence, $\sqrt{t}$ maps to $t^{\frac{p}{2}}=a^{\frac{p-1}{2}}\sqrt{t}$ which determines the value of $Frob_{(t-a),\mathcal{R}^{(2)}}(\sqrt{t})$. Thus, the $(1,2)$-entry of $\operatorname{Frob}_{(t-a),\mathcal{R}^{(2)}}$ is $a^{\frac{p-1}{2}}=\genfrac{(}{)}{}{}{a}{p}$. Similarly, the $(2,1)$-entry of $\operatorname{Frob}_{(t-a),\mathcal{R}^{(2)}}$ is $(1-a)^{\frac{p-1}{2}}=\genfrac{(}{)}{}{}{1-a}{p}$. Combining these results with Proposition 6 and Proposition 7, we obtain the desired result. ∎

### v3

@@ 1. Introduction
Since class field theory was founded in the early 20th century [4], the prime decomposition in abelian extensions of global field has been completely described. In particular, for quadratic extensions, the decomposition pattern is encoded in the Legendre symbol and the desired law corresponds to that of quadratic reciprocity. Combined with Euler’s criterion, the law says
@@
for distinct odd primes $p$ and $q$ such that $4$ divides p1p-1. In the language of class field theory, it relates the Artin symbols at $q$ in two abelian extensions $\mathbb{Q}(\sqrt{p})$ and $\mathbb{Q}(\zeta_{p})$ of $\mathbb{Q}$, where $\zeta_{m}$ denotes a primitive $m$-th root of unity for $m\geq 1$. Despite the success of class field theory in the abelian setting, it remains a fascinating problem to find analogous laws governing the decomposition of primes in nonabelian extensions of global fields.
In this paper, we establish an analogue of (1) for certain nonabelian extensions of function fields. The particular extensions we consider are those arising naturally from attempts to generalize Legendre symbols using Massey products. We mention that Massey products in Galois cohomology have received much attention in their own right, including some recent works [5, 10]. For a specific context regarding the multilinear extension of the Legendre symbol, we refer interested readers to [11] for a survey and to closely related works [3, 9, 1, 2]. Our field extensions first appeared in [6], which is recalled below.
@@
Temporarily, denote by $k$ an arbitrary field containing a primitive $\ell$-th root of unity $\zeta_{\ell}$. In [6], the authors introduced a Galois extension $\mathcal{R}^{(\ell)}$ of the function field $k(t)$ with Galois group $\mathcal{H}(\mathbb{F}_{\ell})$, which we call a mod $\ell$ Heisenberg extension. It is defined as $\mathcal{R}^{(\ell)}\coloneq k(t)(t^{1/\ell},(1-t)^{1/\ell},\epsilon_{\ell}(t)^{1/\ell})$ where $\epsilon_{\ell}(t)\coloneq\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i}t^{1/\ell})^{i}$. They utilized $\mathcal{R}^{(\ell)}$ to establish a relationship between two arithmetic invariants: the mod $\ell$ Milnor invariants $\mu_{\ell}(g;I)$ of Galois elements defined in [8] and the mod $\ell$ Milnor invariants $\mu_{\ell}(J)$ of primes introduced in [11]. Also, they discovered analogies between $\mathcal{R}^{(\ell)}$ and dilogarithmic functions. From this viewpoint, the authors of [12, 14] have explored related structures involving $\ell$-adic Galois polylogarithms and triple $\ell$-th power symbols. We remark that these triple power symbols are regarded as multilinear generalizations of Legendre symbols.
For our purpose, we will take $k=\mathbb{F}_{p}$ for a prime $p$. Since we need a primitive $\ell$-th root of unity, we will always assume $\ell\mid p-1$. Our main result concerns the splitting of the prime ideal (ta)(t-a) for $a\in\mathbb{F}_{p}-\{0,1\}$ in the mod $\ell$ Heisenberg extension $\mathcal{R}^{(\ell)}/\mathbb{F}_{p}(t)$. To state it, we introduce certain polynomials that will play the role of $x^{(p-1)/2}$ for $\mathbb{Q}(\sqrt{p})$ in (1). Put $\epsilon_{\ell,n}(t)=\prod_{i=1}^{\ell-1}(1-\zeta_{\ell}^{i+n}t^{1/\ell})^{i}$. Note that $\epsilon_{\ell,0}(t)$ is literally $\epsilon_{\ell}(t)$ which is a generator of $\mathcal{R}^{(\ell)}$. Define the polynomial
(2) $A_{\ell}(x):=\frac{1}{\ell}\left(\sum_{j=0}^{\ell-1}\epsilon_{\ell,j}(x)^{\frac{p-1}{\ell}}\right).$
For $\ell=2$, which is the case to be treated separately, it is simply
$A_{2}(x)=\frac{1}{2}\left((1-\sqrt{x})^{\frac{p-1}{2}}+(1+\sqrt{x})^{\frac{p-1}{2}}\right).$
Although $A_{2}(x)$ involves $\sqrt{x}$, it is indeed a polynomial in $x$ since it is invariant under the substitution $\sqrt{x}\mapsto-\sqrt{x}$. Similarly, one can show that $A_{\ell}(x)$ is also a polynomial in $x$ using the substitution $t^{1/\ell}\mapsto\zeta_{\ell}t^{1/\ell}$.
@@
In the case $\ell=2$, we have: {restatable}theoremspliteven Let $p$ be an odd prime and $a\in\mathbb{F}_{p}$ such that $a\neq 0,1,\frac{1}{2}$. Let $n_{a}$ be the number of primes above (ta)(t-a) in $\mathcal{R}^{(2)}/\mathbb{F}_{p}(t)$. Then, $n_{a}$ is determined by the value of $A_{2}(a)$;
@@
The reason we separate the case $\ell=2$ is reflected in the group structure of $\mathcal{H}(\mathbb{F}_{\ell})$ which has an element of order $\ell^{2}$ precisely when $\ell=2$. As another feature of the case $\ell=2$, we identify the morphism between algebraic curves underlying $\mathcal{R}^{(\ell)}/k(t)$ which guide the necessary computation. For a general $\ell$, we do not have such geometric interpretation and rely on algebraic manipulations. Despite of the difference, a key ingredient common to all cases is a symmetry with respect to action of the Galois group. The construction of such a basis seems new for $\ell\geq 3$ and it is the most technical part of the paper.
@@ 2.1. Frobenius conjugacy classes
    Let $\mathfrak{q}$ be a prime ideal of $\mathcal{O}_{L}$ lying over $\mathfrak{p}$, and $f$ be the order of $\operatorname{Frob}_{\mathfrak{p},\mathfrak{q}}\in\operatorname{Gal}(L/K)$. Then, the number of prime ideals of $\mathcal{O}_{L}$ lying above $\mathfrak{p}$ equals to $\frac{1}{f}[L:K]$.
@@ 2.3. Mod $\ell$ Heisenberg Extension
Using the notation above, $\mathcal{R}^{(\ell)}/k(t)$ is a Galois extension such that $\operatorname{Gal}(\mathcal{R}^{(\ell)}/k(t))$ is isomorphic to $\mathcal{H}(\mathbb{F}_{\ell})$. Moreover, it is unramified outside $\{\infty,0,1\}$.
@@
induces the isomorphism $\operatorname{Gal}(\mathcal{R}^{(\ell)}/k(t))\cong\mathcal{H}(\mathbb{F}_{\ell}).$
@@ 3. $\ell=2$
Since the Frobenius conjugacy class that corresponds to (ta)(t-a) has order of at most $4$ by Lemma 1, Proposition 1 implies that the number of prime ideals above (ta)(t-a) is at least $2$. ∎
@@
Here, recall that $\mathcal{C}^{(2)}$ is the Fermat plane curve $X^{2}+Y^{2}=Z^{2}$ in $\mathbb{P}^{2}$ that corresponds to $\mathcal{K}^{(2)}$. Since $\genfrac{(}{)}{}{}{a}{p}=\genfrac{(}{)}{}{}{1-a}{p}=1$, the splitting of (ta)(t-a) in $\mathcal{K}^{(2)}$ can be represented by four closed points $(\pm x:\pm y:1)$ on $\mathcal{C}^{(2)}$ where $x,y\in\mathbb{F}_{p}$ that satisfies $x^{2}=a,y^{2}=1-a$.
@@
$\genfrac{(}{)}{}{}{1-x}{p}=\frac{1}{2}\left(\genfrac{(}{)}{}{}{1-x}{p}+\genfrac{(}{)}{}{}{1+x}{p}\right)=\frac{1}{2}\left((1-x)^{\frac{p-1}{2}}+(1+x)^{\frac{p-1}{2}}\right).$
Recall from (2) that $A_{2}(a)=\frac{1}{2}\left((1-\sqrt{a})^{\frac{p-1}{2}}+(1+\sqrt{a})^{\frac{p-1}{2}}\right)$. since $x^{2}=a$ in $\mathbb{F}_{p}$ and $\frac{1}{2}\left((1-x)^{\frac{p-1}{2}}+(1+x)^{\frac{p-1}{2}}\right)$ only has even degree terms, $A_{2}(a)=\genfrac{(}{)}{}{}{1-x}{p}$. Thus, if $\genfrac{(}{)}{}{}{1-x}{p}=1$, the value of the above polynomial is $1$, and if $\genfrac{(}{)}{}{}{1-x}{p}=-1$, the value of the above polynomial is 1-1 and we obtain our desired result. ∎
Now, we show the value of $A_{2}(a)$ determines other quadratic residue symbols.
@@
Note that $A_{2}(a)=x_{\frac{p-1}{2}}$. Then, $x_{n}$ satisfies the recursive relation $x_{n+1}=2x_{n}+(a-1)x_{n-1}$ with initial value $x_{0}=1,x_{1}=1$. Also, from the definition it is straightforward to see that $x_{2n}=2x_{n}^{2}-(1-a)^{n}$ holds. Now, for $a\in\mathbb{F}_{p}$,
$x_{p}=\frac{1}{2}\left((1-\sqrt{a})^{p}+(1+\sqrt{a})^{p}\right)=\frac{1}{2}\sum_{i=0}^{p}(1+(-1)^{i})\binom{p}{i}(\sqrt{a})^{i}=1$
where the first equality is the definition, the second equality is the binomial expansion, and the third equality holds because all but the first term vanishes. In the last equality, we rely on $p$ being odd when $\ell=2$. A similar argument shows that
$x_{p+1}=\frac{1}{2}((1-\sqrt{a})^{p+1}+(1+\sqrt{a})^{p+1})=1+a^{\frac{p+1}{2}}=1+a^{\frac{p-1}{2}}\cdot a=1+\genfrac{(}{)}{}{}{a}{p}a.$
Now, consider the case $\genfrac{(}{)}{}{}{a}{p}=1$. In this case, $x_{p+1}=1+a$ and we get $x_{p-1}=1$ from the recursive formula for $x_{n}$. Then,
$x_{\frac{p-1}{2}}^{2}=\frac{1}{2}\left(1+(1-a)^{\frac{p-1}{2}}\right)=\frac{1}{2}\left(1+\genfrac{(}{)}{}{}{1-a}{p}\right)$.
@@
In the other case $\genfrac{(}{)}{}{}{1-a}{p}=-1$, we have $x_{p+1}=1-a$. Again from the recursive formula, we get $x_{p-1}=\frac{a+1}{1-a}$. Then,
$x_{\frac{p-1}{2}}^{2}=\frac{1}{2}\left(\frac{a+1}{1-a}+(1-a)^{\frac{p-1}{2}}\right)=\frac{1}{2}\left(\frac{1+a}{1-a}+\genfrac{(}{)}{}{}{1-a}{p}\right)$.
If $\genfrac{(}{)}{}{}{1-a}{p}=1$, $x_{\frac{p-1}{2}}^{2}=\frac{1}{1-a}$ which is never $0,1$. If $\genfrac{(}{)}{}{}{1-a}{p}=-1$, $x_{\frac{p-1}{2}}^{2}=\frac{a}{1-a}$ which is not $0,1,\frac{1}{1-a}$ except the case $a=\frac{1}{2}$. Combining these results, we proved the desired result. ∎
@@
From this calculation, we observe that for a representative $(a_{ij})\in\mathcal{H}(\mathbb{F}_{2})$ of a conjugacy class, if $a_{12}=0,a_{23}=1$ or $a_{12}=1,a_{23}=0$ holds, $(a_{ij})$ has order 2. Otherwise, if $a_{12}=a_{23}=1$, $(a_{ij})$ has order 4.
Recall from Proposition 4 that, under the identification
$\operatorname{Gal}(\mathcal{R}^{(2)}/k(t))\cong\mathcal{H}(\mathbb{F}_{2}),$
the $(1,2)$- and $(2,3)$-entries are determined by the images of $\sqrt{t}$ and $\sqrt{1-t}$, respectively. More precisely, for $\sigma\in\operatorname{Gal}(\mathcal{R}^{(2)}/k(t))$, the $(1,2)$-entry of $\sigma$ is given by
$\frac{1}{2}\left(1-\frac{\sigma(\sqrt{t})}{\sqrt{t}}\right),$
while the $(2,3)$-entry is
$\frac{1}{2}\left(1-\frac{\sigma(\sqrt{1-t})}{\sqrt{1-t}}\right).$
Now, we show that each $(1,2),(2,3)$-entry of the Frobenius conjugacy class corresponding to (ta)(t-a) is determined by $\genfrac{(}{)}{}{}{a}{p}$ and $\genfrac{(}{)}{}{}{1-a}{p}$, respectively. First, restriction of $\operatorname{Frob}_{(t-a),\mathcal{R}^{(2)}}$ to $K_{1}=\mathbb{F}_{p}(t)(\sqrt{t})$ is $\operatorname{Frob}_{(t-a),K_{1}}$. Since $K_{1}/\mathbb{F}_{p}(t)$ is abelian, its induced morphism to $\mathcal{O}_{K_{1}}/(t-a)\to\mathcal{O}_{K_{1}}/(t-a)$ is $x\mapsto x^{p}$. Hence, $\sqrt{t}$ maps to $t^{\frac{p}{2}}=a^{\frac{p-1}{2}}\sqrt{t}$ which determines the value of $\operatorname{Frob}_{(t-a),\mathcal{R}^{(2)}}(\sqrt{t})$ as $a^{\frac{p-1}{2}}\sqrt{t}=\genfrac{(}{)}{}{}{a}{p}\sqrt{t}$. Thus, the $(1,2)$-entry of $\operatorname{Frob}_{(t-a),\mathcal{R}^{(2)}}$ is $\frac{1}{2}(1-\genfrac{(}{)}{}{}{a}{p})$. Similarly, the $(2,3)$-entry of $\operatorname{Frob}_{(t-a),\mathcal{R}^{(2)}}$ is $\frac{1}{2}(1-\genfrac{(}{)}{}{}{1-a}{p})$. Combining these results with Proposition 6 and Proposition 7, we obtain the desired result. ∎